# Solving Nonlinear PDEs with Sparse Radial Basis Function Networks

### v1

@@ 2.4. Representer theorem
(2.9) $r_{k}(l_{k})=\begin{cases}\hat{E}(x_{k},l_{k})&\text{for }k\leq K_{1},\\ \hat{B}(x_{k},l_{k})&\text{for }k>K_{1}.\end{cases}$
@@
(2.10) $\min_{\mu\in M(\Omega)}\frac{1}{2}\sum_{k=1}^{K}w_{k}(r_{k}(l_{k}[u_{\mu}]))^{2}+\alpha\lVert\mu\rVert_{M(\Omega)}$
We proceed to prove the representer theorem, which fundamentally relies on variants of the Carathéodory theorem [20, 30]. A key ingredient in the argument is the characterization of the extreme points of certain sets arising in the sparse minimization problem. Recall that an extreme point of a convex set is a point that cannot be expressed as a nontrivial convex combination of two distinct elements in the set. To that end, we first summarize some standard properties of closed balls in $M(\Omega)$ in the following lemma.
Lemma 2.14.
@@
The first statement in the above comes from the Banach-Alaoglu theorem and the lower semicontinuity under the weak-$\ast$ convergence. The second statement is also well-known, see e.g., [4, 9].
@@
(2.11) $\min_{\mu\in M(\Omega)}\|\mu\|_{M(\Omega)}\quad\text{subject to}\quad l[u_{\mu}]=\vec{l}.$
Lemma 2.15.
@@
First of all, by the linearity of the constraint, it is straightforward that $S$ is convex. By the lower semicontinuity of the total variation norm and Lemma 2.4, $S$ is also closed. Finally, since the optimization problem minimizes the total variation norm subject to a constraint, any solution must lie in a norm sublevel set $B_{t}$ for some $t>0$. Hence, $S\subset B_{t}$ is a closed subset of a compact set and is therefore compact. Lastly, by the Krein-Milman theorem [17], $S$ contains at least one extreme point. ∎
The representer theorem is stated below.
Theorem 2.16 (Representer theorem (finite representation)).
@@
According to Theorem 2.5, a (not necessarily finite) solution exists. However, standard representer theorems derived for convex problems are not immediately applicable (see, e.g., [4, Theorem 3.9]), due to the nonlinarity (2.9). We argue by showing that any solution of (2.10) also solves a simpler convex problem. For that we define the optimal linear operator vector as $\bar{l}=l[u_{\bar{\mu}}]\in\mathbb{R}^{\bar{N}}$ It is easy to see that $\bar{\mu}$ is also a minimizer of the following convex problem
$\min\{\lVert\mu\rVert_{M(\Omega)}\;|\;\mu\in M(\Omega);\,l[u_{\mu}]=\bar{l}\},$
and conversely, any solution to this problem is also a solution of ($P_{\mu}^{\text{emp}}$). This problem is convex and admits a solution supported on at most $\bar{N}$ atoms. Specifically, by applying [8, Theorem 3.1], together with Lemmas 2.14 and 2.15, we conclude that there exists a solution expressible as a convex combination of at most $\bar{N}$ extreme points of the norm ball $B_{t^{\ast}}$, where $t^{\ast}$ is the optimal value of ($P_{\mu}^{\text{emp}}$). This yields the desired expression for $\bar{\mu}$ by Lemma 2.14(2). ∎
2.5. Optimality conditions and dual variables
@@
To characterize $\bar{p}$, we require the dual operator of $\mathcal{N}$ and the linear PDE operator in an appropriate sense. However, this can get very technical due to the fact that this dual operator would have to take inputs from the dual space of $C^{2}(\overline{D})$, which is a space of distributions (cf. [60, section 5.2]) To simplify the presentation, we identify $C^{2}(\overline{D})$ with a subspace of $C(\overline{D},\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}^{d^{2}})$, using the embedding
@@
(2.12) $\bar{p}(\omega)=\sum_{k=1}^{K}w_{k}r_{k}(l_{k}[\bar{u}])(\nabla_{l}r_{k}(l_{k}[\bar{u}]))^{T}\widetilde{\varphi}_{k}(\omega)$
@@
Proposition 2.18.
The first order necessary conditions derived in Proposition 2.17 can be expressed with the dual variable (2.12) as the subdifferential inclusion $-\bar{p}\in\alpha\partial\lVert\bar{\mu}\rVert_{M(\Omega)}$ or the support conditions
@@
Proposition 2.19 (Boosting step).
@@
(2.13) $p[u](\omega):=\sum_{k=1}^{K}w_{k}r_{k}(l_{k}[u])(\nabla_{l}r_{k}(l_{k}[u]))^{T}\widetilde{\varphi}_{k}(\omega).$
@@
Remark 2.20.
By direction calculation, the dual variable $p[u](\omega)$ admits the alternative expression
(2.14) $p[u](\omega)=\left.\frac{d}{d\tau}\left[\hat{L}\left(u+\tau\varphi(\cdot;\omega)\right)\right]\right|_{\tau=0},\quad\omega\in\Omega.$
This characterizes $p[u](\omega)$ as the directional derivative of the loss along the candidate feature function $\varphi(\cdot;\omega)$, quantifying the potential decrease in the loss if the node $\omega$ were added. This expression provides an intuitive explanation for Proposition 2.19: the directional derivative must be large enough (in absolute value) to outweigh the corresponding increase in the regularization penalty. This observation forms the cornerstone of our optimization strategy, where kernel nodes are inserted and deleted dynamically.
3. Algorithmic framework
@@
(3.1) $\min_{c,\omega}\hat{L}(u_{c,\omega})+\alpha\|c\|_{1}=\min_{c,\omega}\hat{\ell}(c,\omega)+\alpha\|c\|_{1}$
@@
3.1. Kernel nodes insertion via gradient boosting
@@
According to Proposition 2.19 and Remark 2.20, the dual variable $p[u](\omega)$ provides a good estimate of the potential reduction in the objective function if a new kernel node is added at location $\omega$. In practice, we uniformly sample candidate locations from $\Omega$ and select the one with the largest $|p[u](\omega)|$ that also satisfies $|p[u](\omega)|>\alpha$. We then compare it with the dual variables of existing kernel nodes. The selected node, denoted $\omega_{N+1}$, is inserted into the model $u=\sum_{i=1}^{N}c_{n}\varphi(\cdot;\omega_{n})$ if
(3.2) $|p[u](\omega_{N+1})|>\max_{1\leq n\leq N}|p[u](\omega_{n})|$
@@
3.2. Semi-smooth Gauss-Newton
@@
(3.3) ^$-\nabla\hat{\ell}(c,\omega)=-J_{R}(c,\omega)^{T}WR(c,\omega)\in\alpha\partial_{c,\omega}\|c\|_{1}$
@@
(3.4) $\text{Prox}_{\alpha\|\cdot\|_{1}}(x)&=\text{argmin}_{y}\left[\frac{1}{2}\|x-y\|_{2}^{2}+\alpha\|y\|_{1}\right]\\ &=\left(\text{sign}(x_{i})\max\{|x_{i}|-\alpha,0\}\right)_{i=1}^{n}.\end{aligned}$
@@
(3.5) $G(q,\omega)=(q-\text{Prox}(q),0)+\nabla\hat{\ell}(\text{Prox}(q),\omega)=0.$
$G$ is referred to as the Robinson normal map, introduced in [48]. To address the non-differentiability introduced by the proximal operator, the generalized derivative of $G$ is calculated in a semismooth context [43], given by
(3.6) $DG(q,\omega)=(\text{Id}-DP(q,\omega))+\nabla^{2}\hat{\ell}(\text{Prox}(q),\omega)DP(q,\omega)$
@@
(3.7) $\nabla^{2}\hat{\ell}(c,\omega)\approx J_{R}(c,\omega)^{T}WJ_{R}(c,\omega),$
which is standard in Gauss-Newton type methods [62].
@@
We remark that the optimization problem is in general nonconvex, this results both from when $r_{k}$ is nonlinear in $l_{k}$, as is the case with nonlinear PDEs, and also the optimization in the inner weights. Moreover, the second order optimization algorithm exhibits better efficiency than first order methods. We refer readers to [43] for a complete discussion of semi-smooth calculus involved here, and additional technical details are discussed in Appendix A.
3.3. Computational cost
@@
3.4. Treatment of boundary conditions
@@
Even with an appropriate choice of norm, the penalty-based approach remains suboptimal. Since it relaxes the boundary condition rather than enforcing it exactly, tuning the penalty parameter $\lambda$ becomes critical. While a larger $\lambda$ can improve boundary accuracy and theoretically lead to a better solution, it often slows convergence and may be catastrophic to the obtained solution by overfitting the boundary condition. Considering this, alternative approaches to enforcing Dirichlet boundary conditions beyond the penalty method are of great interest [32, 56]. For the homogeneous Dirichlet case, the solution can be parameterized as
(3.8) $u(x)\approx\gamma(x)\sum_{i=1}^{n}c_{n}\varphi(x;\omega_{n})$
where $\gamma\in C^{2}_{0}(D)$ with $\gamma(x)>0$ for $x\in D$ is a twice continuously differentiable function that vanishes on the boundary $\partial D$. This formulation preserves the theoretical guarantees discussed earlier, provided that $u(x)/\gamma(x)$ satisfies the required regularity. More generally, the method works for any Dirichlet boundary function $f:\partial D\rightarrow\mathbb{R}$ under the assumption that $f$ admits an extension $\bar{f}:D\rightarrow\mathbb{R},\bar{f}\big{|}_{\partial D}=f$ that is twice continuously differentiable. Since such an extension may only exists for smooth domains with $f$ at least twice continuously differentiable, this limits this approach in practice to functions $f$ where such an extension can be easily constructed. We may then set $u(x)=f(x)+\gamma(x)\mathcal{N}(x)$. The extension function $\bar{f}$ and the vanishing function $\gamma$ can often be constructed in low-dimensional domains with simple geometry. However, for complex or high-dimensional geometries, finding such functions becomes significantly more challenging. While there are interesting data-driven approaches, their investigation lies beyond the scope of this work and is left for future research.
3.5. Additional implementation components
@@ 3.5.1. Constraining $\omega$
(3.9) $\sigma(s)=\sigma_{\min}+(\sigma_{\max}-\sigma_{\min})\tau(s),$
@@ 4. Numerical experiments
Table 1. Evaluation Metrics
@@
4.1. Semilinear Poisson equation
@@
$v(x;c,k,R)=\tanh\left(k\left(R_{0}-|x-c|2\right)\right)+1,$
@@
Figure 2. Numerical results using Sparse RBFNet (our) and Gaussian Process (GP) on solving a semilinear Poisson equation: (A) Exact solution; (B) Error of solutions obtained by the GP method with different kernel shape parameters; (C) Convergence of loss and growth of number of kernels ($\alpha=10^{-2}$); (D) Error contours of our method. Each $\star$ represents the location parameters $y_{n}$ of learned kernel nodes, and the dotted circle around which are of radius $\sigma_{n}$ (very few kernel nodes outside of $D$ are omitted).
@@
Table 2. Errors, residual loss, and number of kernels for Sparse RBFNet and GP using grid collocation points. Case with $\alpha=10^{-3}$($10^{-4}$) is solved with solution obtained by $\alpha=10^{-2}$($10^{-3}$) as initialization. Test results are evaluated on a $100\times 100$ grid in $D$. All results are averaged over 10 runs.
@@
We highlight several key advantages of our method for solving equations in high-dimensional spaces compared with other methods. First, by employing a shallow network with smooth feature functions instead of a multi-layer perception (MLP) as is often used in other PINN-based methods, all required derivatives can be computed analytically. This is especially beneficial when computing higher-order derivatives, as it eliminates the need to construct large computational graphs for automatic differentiation or to rely on Monte Carlo-based estimations, thereby reducing both memory usage and computational cost [55]. Second, our method offers greater flexibility compared to GP methods. In a GP formulation, the number of trainable/optimizable parameters scales as $\mathcal{O}(K)$, resulting in substantial computational cost for operations such as Cholesky factorization of a matrix of size $\mathcal{O}(K)\times\mathcal{O}(K)$. Since a large number of collocation points $K$ is typically required to learn complex solutions in high-dimensional spaces, this poses a significant computational bottleneck and requires the use of mini-batch techniques [65, 58]. In contrast, our method inserts new kernel nodes/feature functions only when necessary, thereby keeping the number of trainable (optimizable) parameters manageable, usually much fewer than collocation points. Lastly, the regularization term in our method effectively prevents overfitting, which is particularly important when collocation points are relatively sparse in the domain, as this is often the case in high-dimensional problems.
Table 3. Errors and residual loss of numerical solutions to a 4D semilinear Poisson equation using grid collocation points. Test results are evaluated on a grid of $20^{4}$ in $D\subseteq\mathbb{R}^{4}$. All results are averaged over 10 runs.
@@
4.1.3. Boundary conditions treatments

### v2

@@ 2.4 Representer theorem
$r_{k}(l_{k})=\begin{cases}\hat{E}(x_{k},l_{k})&\text{for }k\leq K_{1},\\ \hat{B}(x_{k},l_{k})&\text{for }k>K_{1}.\end{cases}$ (2.9)
@@
$\min_{\mu\in M(\Omega)}\frac{1}{2}\sum_{k=1}^{K}w_{k}(r_{k}(l_{k}[u_{\mu}]))^{2}+\alpha\lVert\mu\rVert_{M(\Omega)}$ (2.10)
We proceed to prove the representer theorem, which fundamentally relies on variants of the Carathéodory theorem (Dubins, 1962; Klee, 1963). A key ingredient in the argument is the characterization of the extreme points of certain sets arising in the sparse minimization problem. Recall that an extreme point of a convex set is a point that cannot be expressed as a nontrivial convex combination of two distinct elements in the set. To that end, we first summarize some standard properties of closed balls in $M(\Omega)$ in the following lemma.
Lemma 17.
@@
The first statement in the above comes from the Banach-Alaoglu theorem and the lower semicontinuity under the weak-$\ast$ convergence. The second statement is also well-known, see e.g., Bartolucci et al. (2023); Bredies and Carioni (2020).
@@
$\min_{\mu\in M(\Omega)}\|\mu\|_{M(\Omega)}\quad\text{subject to}\quad l[u_{\mu}]=\vec{l}.$ (2.11)
Lemma 18.
@@
First of all, by the linearity of the constraint, it is straightforward that $S$ is convex. By the lower semicontinuity of the total variation norm and Lemma 7, $S$ is also closed. Finally, since the optimization problem minimizes the total variation norm subject to a constraint, any solution must lie in a norm sublevel set $B_{t}$ for some $t>0$. Hence, $S\subset B_{t}$ is a closed subset of a compact set and is therefore compact. Lastly, by the Krein-Milman theorem (Conway, 1994), $S$ contains at least one extreme point. ∎
The representer theorem can now be proved with known arguments  (cf., e.g., Boyer et al., 2019, Section 3.3.2).
Theorem 19 (Representer theorem (finite representation)).
@@
According to Theorem 8, a (not necessarily finite) solution exists. However, standard representer theorems derived for convex problems are not immediately applicable (see, e.g., Bartolucci et al., 2023, Theorem 3.9), due to the nonlinarity (2.9). We argue by showing that any solution of (2.10) also solves a simpler convex problem. For that we define the optimal linear operator vector as $\bar{l}=l[u_{\bar{\mu}}]\in\mathbb{R}^{\bar{N}}$ It is easy to see that $\bar{\mu}$ is also a minimizer of problem (2.11) with data $\vec{l}=\bar{l}$ and conversely, any solution to this problem is also a solution of ($P_{\mu}^{\text{emp}}$). This problem is convex and, by applying (Boyer et al., 2019, Theorem 3.1), together with Lemmas 17 and 18, we conclude that there exists a solution expressible as a convex combination of at most $\bar{N}$ extreme points of the norm ball $B_{t^{\ast}}$, where $t^{\ast}$ is the optimal value of ($P_{\mu}^{\text{emp}}$). This yields the desired expression for $\bar{\mu}$ by Lemma 17(2). ∎
2.5 Optimality conditions and dual variables
@@
To characterize $\bar{p}$, we require the dual operator of $\mathcal{N}$ and the linear PDE operator in an appropriate sense. However, this can get very technical due to the fact that this dual operator would have to take inputs from the dual space of $C^{2}(\overline{D})$, which is a space of distributions (cf. Wachsmuth and Walter, 2024, section 5.2). To simplify the presentation, we identify $C^{2}(\overline{D})$ with a subspace of $C(\overline{D},\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}^{d^{2}})$, using the embedding
@@
$\bar{p}(\omega)=\sum_{k=1}^{K}w_{k}r_{k}(l_{k}[\bar{u}])(\nabla_{l}r_{k}(l_{k}[\bar{u}]))^{T}\widetilde{\varphi}_{k}(\omega)$ (2.12)
@@
Proposition 21.
The first order necessary conditions derived in Proposition 20 can be expressed with the dual variable (2.12) as the subdifferential inclusion $-\bar{p}\in\alpha\partial\lVert\bar{\mu}\rVert_{M(\Omega)}$ or the support conditions
@@
Proposition 22 (Boosting step).
@@
$p[u](\omega):=\sum_{k=1}^{K}w_{k}r_{k}(l_{k}[u])(\nabla_{l}r_{k}(l_{k}[u]))^{T}\widetilde{\varphi}_{k}(\omega).$ (2.13)
@@
Remark 23.
By direct calculation, the dual variable $p[u](\omega)$ admits the alternative expression
$p[u](\omega)=\left.\frac{d}{d\tau}\left[\hat{L}\left(u+\tau\varphi(\cdot;\omega)\right)\right]\right|_{\tau=0},\quad\omega\in\Omega.$ (2.14)
This characterizes $p[u](\omega)$ as the directional derivative of the loss along the candidate feature function $\varphi(\cdot;\omega)$, quantifying the potential decrease in the loss if the node $\omega$ were added. This expression provides an intuitive explanation for Proposition 22: the directional derivative must be large enough (in absolute value) to outweigh the corresponding increase in the regularization penalty. This observation forms the cornerstone of our optimization strategy, where kernel nodes are inserted and deleted dynamically.
3 Algorithmic framework
@@
$\min_{c,\omega}\hat{L}(u_{c,\omega})+\alpha\|c\|_{1}=\min_{c,\omega}\hat{\ell}(c,\omega)+\alpha\|c\|_{1}$ (3.1)
@@
3.1 Kernel nodes insertion via gradient boosting
@@
According to Proposition 22 and Remark 23, the dual variable $p[u](\omega)$ provides a good estimate of the potential reduction in the objective function if a new kernel node is added at location $\omega$. In practice, we uniformly sample candidate locations from $\Omega$ and select the one with the largest $|p[u](\omega)|$ that also satisfies $|p[u](\omega)|>\alpha$. We then compare it with the dual variables of existing kernel nodes. The selected node, denoted $\omega_{N+1}$, is inserted into the model $u=\sum_{i=1}^{N}c_{n}\varphi(\cdot;\omega_{n})$ if
$|p[u](\omega_{N+1})|>\max_{1\leq n\leq N}|p[u](\omega_{n})|$ (3.2)
@@
3.2 Semi-smooth Gauss-Newton
@@
^$-\nabla\hat{\ell}(c,\omega)=-J_{R}(c,\omega)^{T}WR(c,\omega)\in\alpha\partial_{c,\omega}\|c\|_{1}$ (3.3)
@@
$\text{Prox}_{\alpha\|\cdot\|_{1}}(x)&=\text{argmin}_{y}\left[\frac{1}{2}\|x-y\|_{2}^{2}+\alpha\|y\|_{1}\right]\\ &=\left(\text{sign}(x_{i})\max\{|x_{i}|-\alpha,0\}\right)_{i=1}^{n}.\end{aligned}$ (3.4)
@@
$G(q,\omega)=(q-\text{Prox}(q),0)+\nabla\hat{\ell}(\text{Prox}(q),\omega)=0.$ (3.5)
$G$ is referred to as the Robinson normal map, introduced by Robinson (1992). To address the non-differentiability introduced by the proximal operator, the generalized derivative of $G$ is calculated in a semismooth context (Pieper, 2015), given by
$DG(q,\omega)=(\text{Id}-DP(q,\omega))+\nabla^{2}\hat{\ell}(\text{Prox}(q),\omega)DP(q,\omega)$ (3.6)
@@
$\nabla^{2}\hat{\ell}(c,\omega)\approx J_{R}(c,\omega)^{T}WJ_{R}(c,\omega),$ (3.7)
which is standard in Gauss-Newton type methods (Wright, 2006).
@@
We remark that the optimization problem is in general nonconvex, which results both from when $r_{k}$ is nonlinear in $l_{k}$, as is the case with nonlinear PDEs, and also the optimization in the inner weights. Moreover, the second order optimization algorithm exhibits better efficiency than first order methods. We refer readers to Pieper (2015) for a complete discussion of semi-smooth calculus involved here, and additional technical details are discussed in Appendix A.
3.3 Computational cost
@@
3.4 Treatment of boundary conditions
@@
Even with an appropriate choice of norm, the penalty-based approach remains suboptimal. Since it relaxes the boundary condition rather than enforcing it exactly, tuning the penalty parameter $\lambda$ becomes critical. While a larger $\lambda$ can improve boundary accuracy and theoretically lead to a better solution, it often slows convergence and may be catastrophic to the obtained solution by overfitting the boundary condition. Considering this, alternative approaches to enforcing Dirichlet boundary conditions beyond the penalty method are of great interest (Lagaris et al., 1998; Sukumar and Srivastava, 2022). For the homogeneous Dirichlet case, the solution can be parameterized as
$u(x)\approx\gamma(x)\sum_{n=1}^{N}c_{n}\varphi(x;\omega_{n})$ (3.8)
where $\gamma\in C^{2}_{0}(D)$ with $\gamma(x)>0$ for $x\in D$ is a twice continuously differentiable function that vanishes on the boundary $\partial D$. This formulation preserves the theoretical guarantees discussed earlier, provided that $u(x)/\gamma(x)$ satisfies the required regularity. More generally, the method works for any Dirichlet boundary function $f:\partial D\rightarrow\mathbb{R}$ under the assumption that $f$ admits an extension $\bar{f}:D\rightarrow\mathbb{R},\bar{f}\big{|}_{\partial D}=f$ that is twice continuously differentiable. Since such an extension may only exist for smooth domains with $f$ at least twice continuously differentiable, this limits this approach in practice to functions $f$ where such an extension can be easily constructed. We may then set $u(x)=f(x)+\gamma(x)\mathcal{N}(x)$. The extension function $\bar{f}$ and the vanishing function $\gamma$ can often be constructed in low-dimensional domains with simple geometry. However, for complex or high-dimensional geometries, finding such functions becomes significantly more challenging. While there are interesting data-driven approaches, their investigation lies beyond the scope of this work and is left for future research.
3.5 Additional implementation components
@@ 3.5.1 Constraining $\omega$
$\sigma(s)=\sigma_{\min}+(\sigma_{\max}-\sigma_{\min})\tau(s),$ (3.9)
@@ 4 Numerical experiments
Table 1: Evaluation Metrics
@@
4.1 Semilinear Poisson equation
@@
$v(x;c,k,R)=\tanh\left(k\left(R_{0}-|x-c|\right)\right)+1,$
@@
Figure 2: Numerical results using Sparse RBFNet (our) and Gaussian Process (GP) on solving a semilinear Poisson equation: (a) Exact solution; (b) Error of solutions obtained by the GP method with different kernel shape parameters; (c) Convergence of loss and growth of number of kernels ($\alpha=10^{-2}$); (d) Error contours of our method. Each $\star$ represents the location parameter $y_{n}$ of a learned kernel node, and the dotted circle around has radius $\sigma_{n}$ (very few kernel nodes outside of $D$ are omitted).
@@
Table 2: Errors, residual loss, and number of kernels for Sparse RBFNet and GP using grid collocation points. Case with $\alpha=10^{-3}$($10^{-4}$) is solved with solution obtained by $\alpha=10^{-2}$($10^{-3}$) as initialization. Test results are evaluated on a $100\times 100$ grid in $D$. All results are averaged over 10 runs.
@@
We highlight several key advantages of our method for solving equations in high-dimensional spaces compared with other methods. First, by employing a shallow network with smooth feature functions instead of a multi-layer perception (MLP) as is often used in other PINN-based methods, all required derivatives can be computed analytically. This is especially beneficial when computing higher-order derivatives, as it eliminates the need to construct large computational graphs for automatic differentiation or to rely on Monte Carlo-based estimations, thereby reducing both memory usage and computational cost (Sirignano and Spiliopoulos, 2018). Second, our method offers greater flexibility compared to GP methods. In a GP formulation, the number of trainable/optimizable parameters scales as $\mathcal{O}(K)$, resulting in substantial computational cost for operations such as Cholesky factorization of a matrix of size $\mathcal{O}(K)\times\mathcal{O}(K)$. Since a large number of collocation points $K$ is typically required to learn complex solutions in high-dimensional spaces, this poses a significant computational bottleneck and requires the use of mini-batch techniques (Yang and Owhadi, 2023; Tran-Dinh et al., 2020). In contrast, our method inserts new kernel nodes/feature functions only when necessary, thereby keeping the number of trainable (optimizable) parameters manageable, usually much fewer than collocation points. Lastly, the regularization term in our method effectively prevents overfitting, which is particularly important when collocation points are relatively sparse in the domain, as this is often the case in high-dimensional problems.
Table 3: Errors and residual loss of numerical solutions to a 4D semilinear Poisson equation using grid collocation points. Test results are evaluated on a grid of $20^{4}$ in $D\subseteq\mathbb{R}^{4}$. All results are averaged over 10 runs.
@@
4.1.3 Boundary conditions treatments